# RealCustom++: Representing Images as Real-Word for Real-Time Customization

### v1

@@ 6 Conclusion
In this paper, we present a novel customization paradigm, RealCustom++, that, for the first time, represents given subjects as non-conflict real words, thereby disentangling the similarity of given subjects from the controllability of given text by precisely limiting subject influence to relevant parts. This is achieved by gradually customizing the target real word from its general concept to the specific subject within a novel ”train-inference” decoupled framework. Specifically, during training, RealCustom++ learns the general alignment capabilities between visual conditions and all the words in the pre-trained models’ original text conditions. This is achieved through a novel cross-layer cross-scale projector to robustly and finely extract subject features, and a novel curriculum training recipe to adapt the generated subject to diverse poses and sizes. During inference, based on the learned alignment, RealCustom++ customizes the generation of the target real words while keeping other subject-irrelevant regions uncontaminated through a novel adaptive mask guidance. Moreover, we propose a novel multi-real-words customization algorithm to extend RealCustom++ to accommodate multiple-subject customization across various tasks. Extensive experiments demonstrate that RealCustom++ achieves the unity of high-quality similarity and controllability in real-time open-domain scenarios for both single-subject and multiple-subject customization.
Acknowledgments
We would like to thank our colleagues at ByteDance, Yiming Luo, for his valuable help with multi-view data processing. We also thank Wei Liu and Shiqi Sun for their valuable discussions at the very beginning of this research.
@@ Appendix A Evaluation Benchmarks
The full subject images used to evaluate single-subject customization are shown in Figure 18. These images are derived from the standard DreamBench[3] (the first three rows) and are further supplemented with more challenging cases in open domains, such as cartoon characters and buildings, for a more comprehensive evaluation than our conference version. For the first three rows of subject images, the corresponding editing prompts are listed in Table IX, which are also employed from the standard DreamBench benchmark.
editing prompts for non-live objects editing prompts for live objects
”a in the jungle” ”a in the jungle”
”a in the snow” ”a in the snow”
”a on the beach” ”a on the beach”
”a on a cobblestone street” ”a on a cobblestone street”
”a on top of pink fabric” ”a on top of pink fabric”
”a on top of a wooden floor” ”a on top of a wooden floor”
”a with a city in the background” ”a with a city in the background”
”a with a blue house in the background” ”a with a blue house in the background”
”a on top of a purple rug in a forest” ”a on top of a purple rug in a forest”
”a with a wheat field in the background” ”a wearing a red hat”
”a with a tree and autumn leaves in the background” ”a wearing a santa hat”
”a with the Eiffel Tower in the background” ”a wearing rainbow scarf”
”a floating on top of water” ”a wearing a black top hat and a monocle”
”a floating in an ocean of milk” ”a in a checf outfit”
”a on top of green grass with sunflowers around it” ”a in a firefighter outfit”
”a on top of a mirror” ”a in a police outfit”
”a on top of the sidewalk in a crowded street” ”a wearing pink sunglasses”
”a on top of a dirt road” ”a wearing a yellow shirt”
”a on top of a white rug” ”a in a purple wizard outfit”
”a red ” ”a red ”
”a purple ” ”a purple ”
”a shiny ” ”a shiny ”
”a wet ” ”a wet ”
”a cube shaped ” ”a cube shaped ”
TABLE IX: Editing prompt list for quantitative evaluation on single subject customization. For our proposed RealCustom++, “” is replaced with the super-category of the given subject, while for other methods of previous pseudo-word paradigm, “” is replaced with the learned pseudo word, i.e., S*.
Figure 18: The subject images for single-subject customization evaluation, which are derived from the standard DreamBench[3] (the first three row) and are further supplemented with more challenging cases in open domains, such as cartoon characters and buildings, for a more comprehensive evaluation than our conference version.
@@
The subject images and corresponding editing prompts for multiple-subjects customization evaluation are listed in Figure 19.

### v2

@@ 5 Conclusion
In this paper, we present RealCustom++, a novel customization paradigm that represents given subjects as non-conflict real words, disentangling subject similarity from text controllability by precisely limiting subject influence to relevant parts. This is achieved by gradually customizing the target real word from its general concept to the specific subject within a novel ”train-inference” decoupled framework. During training, RealCustom++ learns the general alignment between visual conditions and all words in the pre-trained models’ original text conditions. This is achieved through a novel cross-layer cross-scale projector to robustly extract subject features and a curriculum training recipe to adapt the generated subject to diverse poses and sizes. During inference, based on the learned alignment, RealCustom++ customizes the generation of the target real words while keeping other subject-irrelevant regions uncontaminated through adaptive mask guidance. Additionally, we propose a multi-real-words customization algorithm to extend RealCustom++ for multiple-subject customization across various tasks. Extensive experiments demonstrate that RealCustom++ achieves high-quality similarity and controllability in real-time open-domain scenarios for both single-subject and multiple-subject customization.
@@ Appendix A Evaluation Benchmarks
A.1 Single-subject Customization Evaluation Benchmark
The full subject images used to evaluate single-subject customization are shown in Figure 13. These images are derived from the standard DreamBench[3] (the first three rows) and are further supplemented with more challenging cases in open domains, such as cartoon characters and buildings, for a more comprehensive evaluation than our conference version. For the first three rows of subject images, the corresponding editing prompts are listed in Table VII, which are also employed from the standard DreamBench benchmark.
@@ A.1 Single-subject Customization Evaluation Benchmark
A.2 Multiple-subject Customization Evaluation Benchmark
The subject images and their corresponding editing prompts used for the multiple-subject customization evaluation are listed in Figure 14. This figure provides a comprehensive overview of the input images and the specific textual descriptions used to guide the customization process for each subject.
Appendix B Additional Experiments
We provide more comprehensive ablation experiments.
B.1 Comparison on human ID customization.
Considering that concurrently there are also many methods focusing exclusively on human ID customized generation (a specific limited domain), we further compare our approach with these methods in the human ID scenario to demonstrate our generalization, even though it is not the primary focus of this paper. We consider three recent state-of-the-art methods: PhotoMaker[23], Instant-ID[24], and PuLID[25].
We provide a qualitative comparison with these methods in Figure 15, demonstrating that our open-domain paradigm, RealCustom++, can also effectively generalize to this specific domain, even though human ID customization is not the primary focus of this paper. Specifically, RealCustom++ generates consistent IDs (as shown in the first two rows) while preserving elements such as clothing, hairstyles, and accessories (illustrated in the last three rows), which is challenging for existing methods.
B.2 Visualization of using different attention mask to construct image guidance mask
We visualize their differences in Figure 16, showing that (1) solely relying on cross-attention maps results in scattered mask guidance (e.g., the image guidance mask for the backpack, where holes exist in the center of the mask), causing the degradation of subject details; (2) using all-resolution self-attention maps makes the guidance mask over-focused, resulting in the degradation of mask accuracy.
B.3 Ablation study on early stop regularization
The ablation study on early stop regularization for more temporally stable image mask guidance is provided in Table VIII. This aligns with our motivation, where the image guidance mask tends to converge in the middle diffusion steps, and reusing the middle step’s mask increases both similarity and controllability.
B.4 Visualization of using different Top-K ratios
The qualitative results of using different Top-K ratios $\gamma_{\text{scope}}$ are shown in Figure 17. The visualization results validate our quantitative results in the main paper: (1) Within a proper range (experimentally, $\gamma_{\text{scope}}\in[0.15,0.4]$), the results are quite robust. (2) The maximum normalization in Eq. 15 is important for achieving both high similarity and controllability, as different regions in the selected parts have varying subject relevance and should be assigned different weights. (3) Too small or too large an influence scope will degrade similarity or controllability, respectively.
Figure 17: Visualization of different influence scope.
B.5 Ablation on the impact of using different real words for customization
We further ablate the impact of using different real words for customization in Figure 18. The customized real word is highlighted in red. We show that: (1) In the first row, regardless of how coarse-grained the word is, the customization results of RealCustom++ are quite robust. (2) In the second row, using completely different words to represent the given subject opens a door for a new application: novel concept creation. RealCustom++ combines these two concepts to create a new one, e.g., generating a tiger with the appearance of the given white Bichon Frise, as shown in the second row of Figure 18.
B.6 Ablation on the choice of different feature combinations
We ablate the choice of different feature combinations for the proposed ross-layer cross-scale projector in Table IX. The results validate our hypothesis that same-level features should be added element-wise, while different-level features should be concatenated token-wise to avoid conflicts.
Early Stop Step CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
5 32.06 85.13
10 33.20 87.21
12 33.43 87.32
15 33.40 87.07
20 33.41 86.84
None (w/o early stop regularization) 33.41 86.67
TABLE VIII: Ablation study of mask calculation early stop regularization for more temporally stable image mask guidance.
Feature Combination Setting CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
cross-scale element add + cross-layer token concat 29.94 86.33
cross-scale token concat + cross-layer token concat 33.65 85.62
TABLE IX: Ablation study of different feature combination.
Appendix C Discussion